OpenStudy (yacoub1993):

Solve the logarithmic equation. Be sure to reject any value that is not in the domain of the original logarithmic expressions. Give the exact answer. log6x2 = log6(5x + 36) A. (3/2) B. {9} C. Ø D. {9, -4}

OpenStudy (yacoub1993):

log 6 + logx^2 = log6 + log(5x+36) logx^2 = log(5x+36) x^2 = 5x + 36 lol ....I am stuck after that

OpenStudy (anonymous):

Well I'll continue where you left off: x^2 = 5x + 36 x = 5x(1/2) + 36(1/2) (You have to multiply 1/2 to get rid of the exponent) x = 2.5x + 18 1.5x = 18 x = 12 I think that's right . . .

OpenStudy (anonymous):

\[x ^{2}-5x-36=0\] 1*-36=-36 9-4=5 9*-4=-36 \[x ^{2}-\left( 9-4 \right)x-36=0\] factor and find the values of x,regect any negative value.

OpenStudy (yacoub1993):

@surjithayer the equation u started off with is wrong This is the correct equaton\[x ^{2}-5x+36=0\]

OpenStudy (yacoub1993):

i am not sure though

OpenStudy (anonymous):

\[x ^{2}-9x+4x-36=0\] x(x-9)+4(x-9)=0 (x-9)(x+4)=0 It gives x=9,-4 solution is {-4,9} No need to reject Because(-4)^2=16,which is positive Also 5*-4+36=16 which is positive

OpenStudy (anonymous):

The equation you solved is \[x ^{2}=5x+36\]

OpenStudy (yacoub1993):

ohhh